Construction Engineering Financial

Assignment Description: A contractor has signed a contract to pour the footing, foundation, and slab floor for a building. The contractor uses the costs codes and descriptions from RS Means (see below). The original estimate for the labor to complete the footing and foundation was $8,400 and the contractor has subcontractor the work out for $8,560. The subcontractor has billed the contractor for $2,700 for this work. No additional cost commitments are expected to be made for this line item. The original estimate for concrete to complete the footings and foundation is $9,270, no costs have been committed, and the contractor has received a bill for $3,960. The contractor estimates that it will cost an additional S5,310 to purchase the remaining concrete for the footings and foundations. The original estimate for the labor to complete the slab floor is $20,000 and S100 change order has been approved for additional work. This work will be performed by in-house crews. The work has yet to be performed and the contractor, but the contractor has been given a 1% price discount for the slab concrete from the supplier, which reduces the total estimate for this item by 1%, Determine the total estimated cost at completion for this project and the variance for each cost code 03000 CONCRETE 03200 Under-slab Gravel 03300 Footing & Foundation Labor 03400 Footing & Foundation Concrete 03450 Concrete Pump 03500 Slab/Floor-Labor 03600 Slab/Floor-Concrete 03650 Light-Weight Concrete 03700 Pre-cast Concrete 03900 Rebar
No need to use all rows. Just use what you need and have.
media%2Fe02%2Fe028658b-9df1-4a33-91af-c8
Answer #1
Code Description

Original Estimate

(A)

Changed Orders

(B)

Total Estimate

(C = A+B)

Committed Costs

(D)

Committed Costs Invoiced

(E)

Non Committed Costs Invoiced

(F)

Total Committed and Non Committed Costs Invoiced

(G= E+F)

Estimated Cost to Complete

(H)

Total Estimated Cost at Completion

(I=G+H)

Variance Over/Under

(J=I-C)

03300 Footing and Foundation - Labor $8400 $8400 $8560 $2700 $2700 $2700 -$5700
03400 Footing and Foundation - Concrete $9270 $9270 $3960 $3960 $5310 $9270 $0
03500 Slab/Floor - Labor $20000 $100 $20100

$19899

(Deducting 1% Supplier Discount)

$19899

$19899 -$201
Total $37670 $100 $37770 $28459 $2700 $3960 $6660 $25209 $31869 -$5901